If we consider the effect of taxes, then the degree of operating leverage can be written as: DOL = 1 + [FC × (1 – TC) – TC × D]/OCF Consider a project to supply Detroit with 27,000 tons of machine screws annually for automobile production. You will need an initial $5,300,000 investment in threading equipment to get the project started; the project will last for 6 years. The accounting department estimates that annual fixed costs will be $1,275,000 and that variable costs should be $240 per ton; accounting will depreciate the initial fixed asset investment straight-line to zero over the 6-year project life. It also estimates a salvage value of $650,000 after dismantling costs. The marketing department estimates that the automakers will let the contract at a selling price of $350 per ton. The engineering department estimates you will need an initial net working capital investment of $510,000. You require a return of 14 percent and face a tax rate of 25. a. What is the percentage change in OCF if the units sold changes to 28,000? (Do not round intermediate calculations and enter your answer as a percent rounded to 4 decimal places, e.g., 32.1616.) b. What is the DOL at the base-case level of output? (Do not round intermediate calculations and round your answer to 4 decimal places, e.g., 32.1616.)
Answer #1

1.
=((28000*(350-240)-1275000-5300000/6)*(1-25%)+5300000/6)/((27000*(350-240)-1275000-5300000/6)*(1-25%)+5300000/6)-1=5.52918179279531%

2.
=1+(1275000*(1-25%)-25%*5300000/6)/((28000*(350-240)-1275000-5300000/6)*(1-25%)+5300000/6)=1.46705477639587